One of your friends knows that you are taking a statistics course and asks for your help with her chemistry lab report. She has five measurements on the melting point of a compound: 122.4, 121.8, 122.0. 123.0, and 122.3.
a)The lab manual asks for a 95% confidence interval estimate for the melting point. Show her how to find this estimate.
b) explain to your friend what a 95% confidence interval is.
c)would a 90% confidence interval be narrower than the 95% interval? Explain.

Answer by stanbon(75887)

five measurements on the melting point of a compound:
122.4, 121.8, 122.0. 123.0, and 122.3.
a)The lab manual asks for a 95% confidence interval estimate for the melting point. Show her how to find this estimate.
---
sample mean: 122.3
sample std: 0.4099
standard error = invT(0.975 with df=4)*0.4099/sqrt(5)= 2.7764*0.4099/sqrt(5)
= 0.5090
---
95% CI: 122.3-0.4099 < u < 122.3+0.4099
95% CI: 121.79 < u < 122.71
================================
b) explain to your friend what a 95% confidence interval is.
I'll leave that to you. Try it on some of your friends.
------------------------------------
c)would a 90% confidence interval be narrower than the 95% interval?
Higher confidence requires a broader interval.
===============================================
Cheers,
Stan H..